How much potassium phosphate buffer (1.5M, pH 8) is required to make 100 ml of a 10 mM solution? How much water?


1
Expert's answer
2021-02-01T03:53:08-0500

According to the equation:


C1 × V1 = C2 × V2

where C1 - the initial concentration, V1 - the volume of the initial solution, C2 - the final volume, V2 - the volume of the final solution.


From here, the volume of the initial solution required:

V1 = C2 × V2 / C1 = 10 mM × 100 ml / 1.5 M = 0.01M × 0.1 l / 1.5 M = 6.67 × 10-4 l = 667 ul

As a result, to make the required buffer, take 667 ul of 1.5 M phosphate buffer and 100 ml - 667 ul = 99.3 ml of water